A. Travel agent claims that the average # of rooms in hotels in a large city is 500. At a=0.01, is the claim realistic? The data for a sample of 7 hotels are shown. Give a reason why the claim might be deceptive.
713 300 292 311 598 401 618

Answer by stanbon(75887)

Travel agent claims that the average # of rooms in hotels in a large city is 500. At a=0.01, is the claim realistic? The data for a sample of 7 hotels are shown. Give a reason why the claim might be deceptive.
713 300 292 311 598 401 618
--------
Ho: u = 500 (claim)
Ha: u # 500
-----------------
x-bar = 462 (when rounded up to a whole number)
s(x) = 176
-----
t(462) = (462-500)/[176/sqrt(7)] = -0.5712
-------
p-value = 2*P(t < -0.5712 when df = 6) = 2*tcdf(-100,-0.5712,6) = 0.59
----
Conclusion: Since the p-value is greater than 1%, fail to reject Ho.
---
The test results support the claim.
===================
Cheers,
Stan H.
===================